# Groups with presentations in EDT0L

### 02 Feb 2024 v1

@@ Abstract
To any class of languages LAN, let us associate the class, denoted $\pi(\text{LAN})$, of finitely generated groups that admit a group presentation whose set of relators forms a language in LAN. We show that the class of L-presented groups, as introduced by the first author in 2003, is exactly the class of groups that admit presentations in the class of languages EDT0L. We show that the marked isomorphism problem is not semi-decidable for groups given by EDT0L presentations, contrary to the finite presentation case. We then extend and unify results of the first author with Eick and Hartung about nilpotent and finite quotients, by showing that it is possible to compute the marked hyperbolic and marked metabelian quotients of a group given by an EDT0L presentation. Finally, we show how the results about quotient computations enable the construction of recursively presented groups that do not have EDT0L presentations, thus proving $\pi(\text{EDT0L})\neq\pi(\text{REC})$. This is done by building a residually nilpotent group with solvable word problem whose sequence of maximal nilpotent quotients is non-computable.
@@ 1 Introduction
Such a connection, that has perhaps been little studied, appears via the notion of presentation of groups. For LAN a class of languages, we define a class of groups, which we denote by $\pi(\text{LAN})$, which corresponds to the class of groups that have some presentation in LAN. Namely, it is the class of groups that can be written, for a certain alphabet $S$, as
@@
In fact, presentations of groups are naturally studied in the category of marked groups: a marked group $(G,S)$ is a pair formed by a finitely generated group together with a generating tuple of this group. When a group $G$ is given as a quotient of a free group $\mathbb{F}_{S}$, we naturally get a marking of $G$ by taking as a generating tuple the image of the basis $S$ of the free group in $G$. This leads us to define, associated to a class of languages LAN, a class of marked groups denoted $\pi_{M}(\text{LAN})$, which is the class of marked groups that admit presentations in LAN.
Two classes of groups defined via presentations in classes of languages are well studied: finitely presentable groups and recursively presentable groups. These correspond respectively to $\pi(\text{FIN})$ and $\pi(\text{RE})$, where FIN is the class of finite languages and RE the class of recursively enumerable languages. They are related via Higman’s embedding theorem [Hig61], which says that the set of finitely generated subgroups of finitely presentable groups is exactly the set of recursively presentable groups.
However, most of the theory of formal languages takes place between the classes FIN and RE: FIN and RE could be considered respectively the smallest and biggest classes of languages usually encountered in formal language theory.
Because of this, many presentations that are commonly given as typical examples of recursive presentations in fact provide much more information that an “average” recursive presentation. For example, the lamplighter group $\mathbb{Z}/2\wr\mathbb{Z}$ is given by the presentation
@@
Since “having $H$ as a quotient” is quotient stable exactly when $H$ is the trivial group, recursive presentations do not permit to compute quotients, and in particular the abelianization. In fact, while the set of relators of the presentation of the lamplighter group given above is indeed recursively enumerable, it also belongs to many usually encountered classes of languages that are smaller than RE (including EDT0L).
@@
In view of our previous remarks about formal languages, it seems desirable to relate the notion of L-presentation to a notion of presentation in a certain class of languages. It is rather easy to see that L-presented groups coincide with groups that admit a presentation in DTF0L+FIN (see Section 2.1.2 for a definition), but this class of languages is badly behaved in terms of closure properties, and not considered to be a very interesting class of languages [RS97].
But different classes of languages can yield the same classes of groups: a group has a context-free presentation if and only if it has a finite presentation (Proposition 25), and similarly a group has a recursive set of relators if and only if it has a recursively enumerable set of relators (Craig’s trick). Thus we can search for a class of languages that defines the same set of groups as DTF0L+FIN but which is better behaved.
@@
The class of languages EDT0L is very well behaved in several respects. Recent results of Ciobanu, Dieckert and Elder have opened the way for a wide range of applications of the EDT0L class of languages in group theory, starting with the study of solutions sets to equations in free and hyperbolic groups: it is shown in [CE21] that solution sets to systems of equations and inequations in hyperbolic groups form an EDT0L language whose grammar can be computed in NPSPACE (extending [CDE15] where similar results were proven for free groups). Theorem A is thus the one that really anchors the study of finitely L-presented groups in the theory of formal languages. Moving forwards, it seems that the definition of this class of groups in terms of EDT0L languages should be the preferred one.
@@ 1.2 Marked quotient algorithms
We present here a unified proof of the results of [BEH08] and [Har11], and extend them to the case of metabelian and hyperbolic quotients.
@@
Let $\mathcal{C}$ be any of: the class of nilpotent, of metabelian or of hyperbolic groups. There is an algorithm that, given an EDT0L presentation for a marked group $(G,S)$ and a finite presentation for a marked group $(H,S^{\prime})$ in $\mathcal{C}$, decides whether or not $(H,S^{\prime})$ is a marked quotient of $(G,S)$.
@@ 1.3 Separation of classes of groups defined by languages
One of the classical problems in the theory of formal languages is that of identifying and separating classes of languages: proving that different definitions yield the same class of language, or on the contrary establishing that some inclusions between classes of languages are strict. (The P=NP problem is of this kind, as are all separation problems in complexity theory.) We consider the similar problem for classes of groups defined by presentations. In fact, establishing that $\pi(\text{LAN}_{1})\neq\pi(\text{LAN}_{2})$ is usually much more difficult than establishing the purely language theoretical result $\text{LAN}_{1}\neq\text{LAN}_{2}$. We show why quotient computations permit to establish $\pi(\text{EDT0L})\neq\pi(RE)$.
The results of Theorem C are uniform: for $\mathcal{C}$ one of the classes of nilpotent, metabelian or hyperbolic groups, there is an algorithm which, given as input both an EDT0L presentation for a marked group $(G,S)$ and a finite presentation for a marked group in $\mathcal{C}$, determines whether or not the latter is a marked quotient of the former. We can consider a weaker statement, the non-uniform version of the above result: “for each group $G$ that admits an EDT0L presentation, there is an algorithm that, given a finite presentation for a marked group in $\mathcal{C}$, determines whether it is a marked quotient of $G$”. The point here is that this non-uniform version of Theorem C provides a possibly non-trivial property of each group that admits an EDT0L presentation. In fact, if we take $\mathcal{C}$ to be the class of metabelian groups or the class of free groups, the obtained property is trivial: for each finitely generated group, there is an algorithm that recognizes its marked metabelian quotients, and one that determines its marked free quotients. Thus the interest of Theorem C in regards to free and metabelian quotients is purely in its uniformity.
@@
(The above result obviously remains true replacing “finite groups” by “hyperbolic groups”.) In view of Theorem C, the group given in this theorem cannot have a presentation in EDT0L, even though it is recursively presented. Thus we know that the clauses of Theorem C that concern metabelian groups and free groups do not provide information on groups with presentations in EDT0L, while the one that concerns hyperbolic groups does. The last remaining case is that of nilpotent quotients. We prove here:
@@ 1.4 Open problems
Moving forward, several directions of research appear promising. First, establishing for classes of languages other than EDT0L whether they give rise to groups that admit marked quotient algorithms. Secondly, finding methods other than the construction of groups with non-computable quotients to separate classes of groups defined by languages. For instance we ask:
@@
Find a class of languages LAN such that
@@ 2.1 Language Theory
Let $A$ be any set, to which we refer from now on as the alphabet. The free monoid over $A$, i.e. the set of all words over $A$, including the empty word which we denote by $\epsilon$, is denoted by $A^{*}$. If not specified, the term morphism always refers to monoid homomorphism. A formal language over $A$ is a set of words over $A$, i.e. a subset of $A^{*}$. A class of languages over $A$ is a set of formal languages over $A$.
A class of languages is then defined to be a map that associates to a set $A$ a class of languages over $A$.
In practice (see [RS97]), classes of languages come equipped with ways of giving finite descriptions of the considered languages, that take the form of machines (automata, Turing machines, and so on) or of grammars. We could thus give a more precise definition of the notion of “class of languages”, that reflects more precisely how this term is used, thanks to the theory of numberings, see for instance [Wei87]. But we will not use such a definition here.
We will usually denote classes of languages in all capitals, such as LAN.
@@ 2.1.1 Regular and context free languages
Here, we introduce regular and context free languages, denoted REG and CF. Denote FIN the class of finite languages. In Section 2.2.2, we show that in fact:
@@ 2.1.2 Some L-languages
This section will introduce concepts from the theory of parallel rewriting language classes, which are often called L-languages or “Lindenmayer languages”. For the interested reader, a broader account can be found in chapter 5 of [RS97].
@@ A HDT0L system $(w_{0},\Phi,f)$ defines a language
We adopt the common terminology of saying that a formal language is DT0L (DTF0L, EDT0L,…) if there is a corresponding system which defines that language. We also denote by DT0L, DTF0L, … the respective classes of languages. Additionally, FIN denotes the class of all finite languages.
@@
The class DTF0L trivially contains all finite languages. The words of the finite language can be taken as the seed words of the DTF0L system and the morphisms to be the identity.
@@
The classes EDT0L and HDT0L are effectively the same, i.e. there is an algorithm that computes from an EDT0L system an HDT0L system that defines the same language and vice versa.
@@
The class EDT0L is effectively closed under finite union, applying morphisms and intersection with a regular language.
We denote by DTF0L+FIN the class of languages that are a union of a DTF0L and a finite language.
@@
For being a subclass it suffices, due to the closure properties of EDT0L, to show that both DTF0L and FIN are subclasses of EDT0L. This is achieved by two very similar constructions: One defines a single non-terminal $N$ which is also the seed of the system. Additionally, one adds morphisms that map $N$ to each word in the finite language or seed word of the DTF0L system respectively and that act trivially everywhere else. In the DTF0L case, the morphisms of the original system are added as well, acting trivially on $N$.
@@ 2.1.3 Regular Control
The class of languages defined by regularly controlled EDT0L systems is effectively identical to the class of EDT0L languages.
@@
If one removes from Definition 16 the intersection with terminal words $A^{*}$, so in fact requires the regular control $\mathcal{R}$ to consist only of morphisms mapping $w_{0}$ to $A^{*}$, the defined class of languages is still the same.
@@
2.2 Presentations of groups in classes of languages
@@ 2.2.1 Marked groups and presentations
Let $k$ be natural number. A $k$-marked group is a finitely generated group $G$ together with a $k$-tuple $S=(s_{1},...,s_{k})$ of elements of $G$ that generate it. Note that repetitions are allowed in $S$, the order of the elements matters, and $S$ could contain the identity element of $G$. A morphism of marked groups between $k$-marked groups $(G,(s_{1},...,s_{k}))$ and $(H,(t_{1},...,t_{k}))$ is a group morphism $\varphi$ between $G$ and $H$ that additionally satisfies $\varphi(s_{i})=t_{i}$, $i\in\{1,..,k\}$. It is an isomorphism of marked groups if $\varphi$ is a group isomorphism. Marked groups are considered up to isomorphism.
@@
Let $(G,S)$ be a marked group and LAN a class of languages. We say that $(G,S)$ has a LAN-presentation if $G$ has a group presentation
@@
If LAN is a class of languages, define:
@@
Note that an important point which is not apparent in these definitions is the following: in practice, classes of languages come equipped with a type of finite description for the languages they contain (usually in the form of a grammar). And thus the definition above also allows to define a type of finite description for marked groups: a tuple $S$ of symbols for generators, together with the finite description of a language over $S\cup S^{-1}$, will constitute a finite description of a marked group. It is these finite descriptions that permit us to ask decidability questions about groups given by presentations.
@@ 2.2.2 First examples for the non-injectivity of $\pi$ and $\pi_{M}$
Some of the smallest classes of languages usually considered are finite and regular languages, classes denoted FIN and REG. On the other hand of the spectrum, the biggest classes usually considered are recursive (REC) and recursively enumerable (RE) languages. In each case these different classes of languages yield the same classes of (marked) groups.
@@
Let CF be the class of context free languages. Then $\pi_{M}(\text{CF})=\pi_{M}(\text{FIN})$

### 09 Apr 2024 v2

@@ Abstract
To any family of languages LAN, let us associate the class, denoted $\pi(\text{LAN})$, of finitely generated groups that admit a group presentation whose set of relators forms a language in LAN. We show that the class of L-presented groups, as introduced by the first author in 2003, is exactly the class of groups that admit presentations in the family of languages EDT0L. We show that the marked isomorphism problem is not semi-decidable for groups given by EDT0L presentations, contrary to the finite presentation case. We then extend and unify results of the first author with Eick and Hartung about nilpotent and finite quotients, by showing that it is possible to compute the marked hyperbolic and marked abelian-by-nilpotent quotients of a group given by an EDT0L presentation. Finally, we show how the results about quotient computations enable the construction of recursively presented groups that do not have EDT0L presentations, thus proving $\pi(\text{EDT0L})\neq\pi(\text{REC})$. This is done by building a residually nilpotent group with solvable word problem whose sequence of maximal nilpotent quotients is non-computable.
@@ 1 Introduction
Such a connection, that has perhaps been little studied, appears via the notion of presentation of groups. For LAN a family of languages, we define a class of groups, which we denote by $\pi(\text{LAN})$, which corresponds to the class of groups that have some presentation in LAN. Namely, it is the class of groups that can be written, for a certain alphabet $S$, as
@@
In fact, presentations of groups are naturally studied in the category of marked groups: a marked group $(G,S)$ is a pair formed by a finitely generated group together with a generating tuple of this group. When a group $G$ is given as a quotient of a free group $\mathbb{F}_{S}$, we naturally get a marking of $G$ by taking as a generating tuple the image of the basis $S$ of the free group in $G$. This leads us to define, associated to a family of languages LAN, a class of marked groups denoted $\pi_{M}(\text{LAN})$, which is the class of marked groups that admit presentations in LAN.
Two classes of groups defined via presentations in families of languages are well studied: finitely presentable groups and recursively presentable groups. These correspond respectively to $\pi(\text{FIN})$ and $\pi(\text{RE})$, where FIN is the family of finite languages and RE the family of recursively enumerable languages. They are related via Higman’s embedding theorem [Hig61], which says that the set of finitely generated subgroups of finitely presentable groups is exactly the set of recursively presentable groups.
However, most of the theory of formal languages takes place between the families FIN and RE: FIN and RE could be considered respectively the smallest and biggest families of languages usually encountered in formal language theory.
Because of this, many presentations that are commonly given as typical examples of recursive presentations in fact provide much more information than an “average” recursive presentation. For example, the lamplighter group $\mathbb{Z}/2\wr\mathbb{Z}$ is given by the presentation
@@
Since “having $H$ as a quotient” is quotient stable exactly when $H$ is the trivial group, recursive presentations do not permit to compute quotients, and in particular the abelianization. In fact, while the set of relators of the presentation of the lamplighter group given above is indeed recursively enumerable, it also belongs to many usually encountered families of languages that are smaller than RE (including EDT0L).
@@
In view of our previous remarks about formal languages, it seems desirable to relate the notion of L-presentation to a notion of presentation in a certain family of languages. It is rather easy to see that L-presented groups coincide with groups that admit a presentation in DTF0L+FIN (see Section 2.1.2 for a definition), but this family of languages is badly behaved in terms of closure properties, and not considered to be a very interesting family of languages [RS97].
But different families of languages can yield the same classes of groups: a group has a context-free presentation if and only if it has a finite presentation (Proposition 25), and similarly a group has a recursive set of relators if and only if it has a recursively enumerable set of relators (Craig’s trick). Thus we can search for a family of languages that defines the same set of groups as DTF0L+FIN but which is better behaved.
@@
The family of languages EDT0L is very well behaved in several respects. Recent results of Ciobanu, Dieckert and Elder have opened the way for a wide range of applications of the EDT0L family of languages in group theory, starting with the study of solutions sets to equations in free and hyperbolic groups: it is shown in [CE21] that solution sets to systems of equations and inequations in hyperbolic groups form an EDT0L language whose grammar can be computed in NPSPACE (extending [CDE15] where similar results were proven for free groups). Theorem A is thus the one that really anchors the study of finitely L-presented groups in the theory of formal languages. Moving forwards, it seems that the definition of this class of groups in terms of EDT0L languages should be the preferred one.
@@ 1.2 Marked quotient algorithms
We present here a unified proof of the results of [BEH08] and [Har11], and extend them to the case of abelian-by-nilpotent, hyperbolic and virtual direct product of hyperbolic groups. What we in fact show is that the hypotheses of Noetherianity and of uniform solvability of the word problem, that were used in the case of finite and nilpotent groups, can be replaced more generally by hypotheses of equational Noetherianity and of solvability of the universal Horn theory. Solvability of the universal theory of virtual direct products of hyperbolic groups was established in [CHR20], building on [DG10].
@@
Let $\mathcal{C}$ be any of: the class of abelian-by-nilpotent groups or the class of virtual direct products of finitely many hyperbolic groups. There is an algorithm that, given an EDT0L presentation for a marked group $(G,S)$ and a finite presentation for a marked group $(H,S^{\prime})$ in $\mathcal{C}$, decides whether or not $(H,S^{\prime})$ is a marked quotient of $(G,S)$.
@@ 1.3 Separation of classes of groups defined by languages
One of the classical problems in the theory of formal languages is that of identifying and separating families of languages: proving that different definitions yield the same family of language, or on the contrary establishing that some inclusions between families of languages are strict. (The P=NP problem is of this kind, as are all separation problems in complexity theory.) We consider the similar problem for classes of groups defined by presentations. In fact, establishing that $\pi(\text{LAN}_{1})\neq\pi(\text{LAN}_{2})$ is usually much more difficult than establishing the purely language theoretical result $\text{LAN}_{1}\neq\text{LAN}_{2}$. We show why quotient computations permit to establish $\pi(\text{EDT0L})\neq\pi(RE)$.
The results of Theorem C are uniform: for $\mathcal{C}$ one of the classes of abelian-by-nilpotent or of virtual direct product of hyperbolic groups, there is an algorithm which, given as input both an EDT0L presentation for a marked group $(G,S)$ and a finite presentation for a marked group in $\mathcal{C}$, determines whether or not the latter is a marked quotient of the former. We can consider a weaker statement, the non-uniform version of the above result: “for each group $G$ that admits an EDT0L presentation, there is an algorithm that, given a finite presentation for a marked group in $\mathcal{C}$, determines whether it is a marked quotient of $G$”. The point here is that this non-uniform version of Theorem C provides a possibly non-trivial property of each group that admits an EDT0L presentation. In fact, if we take $\mathcal{C}$ to be the class of abelian-by-$k$-nilpotent groups (for a fixed $k$) or the class of free groups, the obtained property is trivial: for each finitely generated group, there is an algorithm that recognizes its marked abelian-by-$k$-nilpotent quotients, and one that determines its marked free quotients. Thus the interest of Theorem C in regards to free and abelian-by-$k$-nilpotent quotients is purely in its uniformity.
@@
(The above result obviously remains true replacing “finite groups” by “hyperbolic groups”.) In view of Theorem C, the group given in this theorem cannot have a presentation in EDT0L, even though it is recursively presented. Thus we know that the clauses of Theorem C that concern abelian-by-$k$-nilpotent groups and free groups do not provide information on groups with presentations in EDT0L, while the one that concerns hyperbolic groups does. We finally consider nilpotent quotients, without fixing a bound on the nilpotency class. We prove here:
@@ 1.4 Open problems
Moving forward, several directions of research appear promising. First, establishing for families of languages other than EDT0L whether they give rise to groups that admit marked quotient algorithms. Secondly, finding methods other than the construction of groups with non-computable quotients to separate classes of groups defined by languages. For instance we ask:
@@
Find a family of languages LAN such that
@@ 2.1 Language Theory
Let $A$ be any set, to which we refer from now on as the alphabet. The free monoid over $A$, i.e. the set of all words over $A$, including the empty word which we denote by $\epsilon$, is denoted by $A^{*}$. If not specified, the term morphism always refers to monoid homomorphism. A formal language over $A$ is a set of words over $A$, i.e. a subset of $A^{*}$. A family of languages over $A$ is a set of formal languages over $A$.
A family of languages is then defined to be a map that associates to a set $A$ a family of languages over $A$.
In practice (see [RS97]), families of languages come equipped with ways of giving finite descriptions of the considered languages, that take the form of machines (automata, Turing machines, and so on) or of grammars. We could thus give a more precise definition of the notion of “family of languages”, that reflects more precisely how this term is used, thanks to the theory of numberings, see for instance [Wei87]. But we will not use such a definition here.
We will usually denote families of languages in all capitals, such as LAN.
@@ 2.1.1 Regular and context free languages
Here, we introduce regular and context free languages, denoted REG and CF. Denote by FIN the family of finite languages. In Section 2.2.2, we show that in fact:
@@ 2.1.2 Some L-languages
This section will introduce concepts from the theory of parallel rewriting language families, which are often called L-languages or “Lindenmayer languages”. For the interested reader, a broader account can be found in chapter 5 of [RS97].
@@ A HDT0L system $(w_{0},\Phi,f)$ defines a language
We adopt the common terminology of saying that a formal language is DT0L (DTF0L, EDT0L,…) if there is a corresponding system which defines that language. We also denote by DT0L, DTF0L, … the respective families of languages. Additionally, FIN denotes the family of all finite languages.
@@
The family DTF0L trivially contains all finite languages. The words of the finite language can be taken as the seed words of the DTF0L system and the morphisms to be the identity.
@@
The families EDT0L and HDT0L are effectively the same, i.e. there is an algorithm that computes from an EDT0L system an HDT0L system that defines the same language and vice versa.
@@
The family EDT0L is effectively closed under finite union, applying morphisms and intersection with a regular language.
We denote by DTF0L+FIN the family of languages that are a union of a DTF0L and a finite language.
@@
For being a subfamily it suffices, due to the closure properties of EDT0L, to show that both DTF0L and FIN are subfamilies of EDT0L. This is achieved by two very similar constructions: One defines a single non-terminal $N$ which is also the seed of the system. Additionally, one adds morphisms that map $N$ to each word in the finite language or seed word of the DTF0L system respectively and that act trivially everywhere else. In the DTF0L case, the morphisms of the original system are added as well, acting trivially on $N$.
@@ 2.1.3 Regular Control
The family of languages defined by regularly controlled EDT0L systems is effectively identical to the family of EDT0L languages.
@@
If one removes from Definition 16 the intersection with terminal words $A^{*}$, so in fact requires the regular control $\mathcal{R}$ to consist only of morphisms mapping $w_{0}$ to $A^{*}$, the defined family of languages is still the same.
@@
2.2 Presentations of groups in families of languages
@@ 2.2.1 Marked groups and presentations
Let $k$ be a natural number. A $k$-marked group is a finitely generated group $G$ together with a $k$-tuple $S=(s_{1},...,s_{k})$ of elements of $G$ that generate it. Note that repetitions are allowed in $S$, the order of the elements matters, and $S$ could contain the identity element of $G$. A morphism of marked groups between $k$-marked groups $(G,(s_{1},...,s_{k}))$ and $(H,(t_{1},...,t_{k}))$ is a group morphism $\varphi$ between $G$ and $H$ that additionally satisfies $\varphi(s_{i})=t_{i}$, $i\in\{1,..,k\}$. It is an isomorphism of marked groups if $\varphi$ is a group isomorphism. Marked groups are considered up to isomorphism.
@@
Let $(G,S)$ be a marked group and LAN a family of languages. We say that $(G,S)$ has a LAN-presentation if $G$ has a group presentation
@@
If LAN is a family of languages, define:
@@
Note that an important point which is not apparent in these definitions is the following: in practice, families of languages come equipped with a type of finite description for the languages they contain (usually in the form of a grammar). And thus the definition above also allows to define a type of finite description for marked groups: a tuple $S$ of symbols for generators, together with the finite description of a language over $S\cup S^{-1}$, will constitute a finite description of a marked group. It is these finite descriptions that permit us to ask decidability questions about groups given by presentations.
@@ 2.2.2 First examples for the non-injectivity of $\pi$ and $\pi_{M}$
Some of the smallest families of languages usually considered are finite and regular languages, families denoted FIN and REG. On the other hand of the spectrum, the biggest families usually considered are recursive (REC) and recursively enumerable (RE) languages. In each case these different families of languages yield the same classes of (marked) groups.
@@
Let CF be the family of context free languages. Then $\pi_{M}(\text{CF})=\pi_{M}(\text{FIN})$